how to solve logbase7 (x+5)-logbase7 *3=logbase7 *21
Answer by rothauserc(4718)

logbase7 (x+5)-logbase7 *3=logbase7 *21
I assume you meant
logbase7 (x+5)-logbase7 (3) = logbase7 (21)
logbase7 (x+5) = logbase7 (21) + logbase7 (3)
logbase7 (x+5) = 1.565 + 0.565 = 2.13
use definition of logarithm, x=b^y then y=log base b (x)
x+5 = 7^(2.13) = 63.104
x = 58.104